OpenStudy (pokemon23):

Solve each equation for x: 5-2(3x-4)=-11 Solve each equation for x: 5-2(3x-4)=-11 @Mathematics

OpenStudy (pokemon23):

Saifoo :)

OpenStudy (pokemon23):

4x+9=2x+31

OpenStudy (saifoo.khan):

-2(3x-4) = -11 -5 -2(3x-4) = -16 3x - 4 = 8 3x = 8 + 4 x = 12/3 x = 4

OpenStudy (saifoo.khan):

what's that?

OpenStudy (pokemon23):

another math question